OpenStudy (anonymous):

what is the slope of the line with equation 3x+4y=7

OpenStudy (saifoo.khan):

Solve for y.

OpenStudy (saifoo.khan):

y = mx + b where m is the slope

OpenStudy (aravindg):

write this as y=mx+c where m is slope

OpenStudy (aravindg):

here y=-3/4x+7/4 so slope =-3/4

OpenStudy (aravindg):

gt it?

OpenStudy (anonymous):

yea thanks guys